# Benchmarking ERP Analysis: Manual Features, Deep Learning, and Foundation Models

### v1

@@ 4.2.10 EEGConformer
4.3 Foundation Models
4.3.1 BIOT
4.3.2 LaBraM
LaBraM [17] is a foundation model for learning universal EEG representations through large-scale unsupervised pre-training. The framework’s core lies in a vector-quantization neural spectral prediction method: First, a neural tokenizer is trained to encode continuous EEG segments into a compact neural lexicon by reconstructing the Fourier spectrum of the raw EEG signals. Subsequently, a neural Transformer is pre-trained using masked EEG modeling (randomly masking portions of EEG segments and predicting their corresponding neural lexical entries) based on this lexicon. Equipped with learnable temporal and spatial embeddings, this model adaptively learns from EEG inputs of arbitrary channel counts and durations. This method uses more than 2,000 hours of EEG data for pre-training. We load their pre-trained weights for fine-tuning.
4.3.3 CBraMod
4.4 Patch Embedding Comparison for ERP Transformer
Existing EEG Transformer models commonly employ three types of token embeddings: multi-variate, uni-variate, and whole-variate. For example, Medformer [74] employs multivariate patch embedding, PatchTST [72] adopts univariate patch embedding, and iTransformer [73] uses whole-variate patch embedding. Given a multivariate EEG input sample and a patch length , multi-variate embedding uses patches , univariate embedding uses patches , and whole-variate embedding uses patches . These patches are then linearly projected into token embeddings that serve as inputs to the Transformer encoder. Figure 5 illustrates the three embedding strategies. To investigate which embedding method is most suitable for ERP data classification, we implement a Transformer model and evaluate three token-embedding designs. All other components of the Transformer architecture are kept identical across settings, allowing a controlled comparison of the embedding methods. We hope this comparison could inspire future works on designing Transformers for ERP analysis.
5 Experiments
5.1 Experimental Setups
5.2 ERP Stimulus Classification Results
We first investigate stimulus-type classification in ERP data, with results summarized in Table 2. All experiments are conducted under a subject-independent evaluation setup, aiming to identify stimulus-related ERP patterns that generalize across subjects. The Top-1, Top-2, and Top-3 performing methods are highlighted in red, blue, and green, respectively. Across the six datasets (CESCA-AOOD, CESCA-VOOD, CESCA-FLANKER, mTBI-ODD, NSERP-MSIT, and NSERP-ODD), the best method achieved F1 scores of 57.18%, 69.64%, 64.33%, 65.79%, 38.51%, and 68.40%, respectively. These results are substantially higher than random performance, indicating that meaningful and consistent stimulus-related ERP patterns can be extracted even under cross-subject evaluation. Among the three CESCA datasets, the visual oddball and flanker paradigms exhibit higher discriminability than the auditory oddball paradigm. Since these datasets share the same group of subjects and recording conditions, this observation suggests that visual oddball and flanker stimuli elicit more salient and distinguishable neural responses than passive auditory oddball stimuli. For the comparison of method performance, we observe that the top-tier results are predominantly achieved by deep learning models trained from scratch. In contrast, existing foundation models for medical time series and EEG do not demonstrate clear performance advantages when fine-tuned from their released checkpoints. Moreover, manually extracted features rarely achieve top-tier performance, whereas ERP features consistently outperform EEG features in this task.
@@ 5.3 ERP-Based Disease Detection Results
We then investigate ERP-based disease detection performance, with results summarized in Table 3. All experiments are conducted under a subject-independent evaluation setup, aiming to identify disease-specific ERP features that generalize across subjects for real-world disease detection applications. Across the six datasets (PD-SIM, PD-ODD, ADHD-WMRI, SCPD, RLPD, and AOPD), the best method achieves F1 scores of 66.79%, 67.37%, 66.43%, 74.33%, 73.87%, and 69.26%, respectively. Overall, no single ERP paradigm consistently demonstrates a clear advantage for disease detection. Notably, the three smaller Parkinson’s disease datasets, SCPD, RLPD, and AOPD, achieve relatively higher performance than the two larger datasets, PD-SIM and PD-ODD. This trend may be attributed to the more balanced data distributions in the smaller datasets, which contain exactly equal numbers of Parkinson’s disease and healthy control subjects. For method comparison, top-tier performance is primarily achieved by deep learning models trained from scratch and by models fine-tuned from foundation model checkpoints. Two methods based on manually extracted features never achieve top-tier results. Besides, unlike ERP stimulus-type classification, ERP-specific features do not consistently outperform generic EEG features in disease detection. In several datasets, ERP-specific features even perform worse, suggesting that the disease-related information in these datasets may be less tightly aligned with canonical ERP components.
@@ 5.4 Overall Performance Analysis
To provide a comprehensive comparison of different methods across multiple tasks and evaluation metrics, we further analyze the overall performance of 15 methods over 12 datasets 3 evaluation metrics, resulting in 36 evaluations in total. Figure 4 presents the average rank of each method across all 36 evaluations, along with a heatmap showing the individual ranks for each method on each evaluation. Lower ranks and a deeper blue color indicate better performance.
@@
We observe that supervised deep learning models trained from scratch consistently outperform other methods in overall performance. Among all evaluated approaches, EEGConformer achieves the lowest average rank of 3.96, indicating the most stable and competitive performance across diverse tasks and evaluation metrics. Following EEGConformer, ModernTCN, TimesNet, CBraMod, and Medformer constitute the top-tier group. Notably, only CBraMod belongs to the foundation model category, while the remaining methods are trained from scratch. This suggests that existing foundation models can partially transfer to ERP classification tasks. One possible explanation is that current EEG foundation models are rarely pre-trained on ERP-specific data. Most pre-training resources consist of long-sequence spontaneous EEG and do not incorporate ERP-specific preprocessing steps, such as baseline correction or event-based trial epoching. As a result, the representations learned during pre-training may not be optimally aligned with ERP-centric downstream tasks. Manual feature-based methods consistently perform the worst overall. Between the two handcrafted feature methods, ERP-specific features perform relatively better than generic EEG features. These results highlight the limitations of handcrafted statistical and frequency-domain features, which struggle to maintain robust performance across diverse ERP paradigms and disease conditions, particularly under subject-independent evaluation. This suggests that, beyond interpretability or neuroscience-oriented analysis, manual feature extraction offers limited advantages in real-world applications where end-to-end performance is the primary objective.
| Multi-Variate (0.864M) | Uni-Variate (0.818M) | Whole-Variate (0.822M) | |
| CESCA-AODD | 53.260.70 | 54.020.73 | 53.350.81 |
| CESCA-VODD | 67.001.62 | 67.711.77 | 65.661.33 |
| CESCA-FLANKER | 63.210.84 | 63.810.82 | 63.351.39 |
| mTBI-ODD | 63.411.62 | 64.511.83 | 63.981.67 |
| NSERP-MSIT | 37.342.21 | 36.872.85 | 35.692.37 |
| NSERP-ODD | 65.362.32 | 63.592.61 | 62.392.69 |
| PD-SIM | 58.165.05 | 68.411.73 | 67.142.55 |
| PD-ODD | 62.453.19 | 68.500.44 | 66.131.58 |
| ADHD-WMRI | 63.775.06 | 60.533.15 | 60.843.37 |
| SCPD | 67.846.87 | 67.915.04 | 65.445.03 |
| RLPD | 64.474.74 | 60.734.92 | 59.485.00 |
| AOPD | 65.296.78 | 59.436.76 | 59.196.45 |
5.5 Patch Embedding Comparison for Transformer
@@
We compare three commonly used token embedding methods in existing EEG Transformer models to examine their suitability for ERP data classification. All other components, including the self-attention modules, feedforward networks, and layer normalization layers, are kept identical across methods. To ensure comparable model capacities, the patch length is set to 25 for the multi-variate and 100 for the uni-variate embedding methods. As a result, the total number of parameters is 0.864, 0.818, and 0.822 million for the multi-variate, uni-variate, and whole-variate token embedding methods, respectively.
@@
All results are reported in terms of F1 score in Table 4. The uni-variate embedding method achieves the best performance on 7 out of 12 datasets, while the multi-variate embedding method achieves the best results on the remaining 5 datasets. In contrast, the whole-variate embedding method does not achieve top performance on any dataset. These results demonstrate the advantage of uni-variate token embedding for ERP classification and suggest promising directions for designing ERP-specific Transformer backbones.
@@ 6 Discussion
Overall, performance on both stimulus-type classification and brain disease detection remains relatively limited for real-world deployment. This limitation may stem from a combination of dataset characteristics and methodological choices. Although the number of trials in existing ERP datasets is often sufficient for traditional manual feature-based analysis, it may still be inadequate for training data-hungry deep learning models. In addition, dataset quality factors, such as noise contamination and class imbalance, can further constrain achievable performance, which is consistent with our observations in Table 3.
From a methodological perspective, EEGConformer, a model proposed several years ago, achieves the best average ranking in overall performance, outperforming more recent approaches, including foundation models. This observation suggests that recent advances in general EEG representation learning do not directly translate to the ERP sub-domain, highlighting the need for backbone architectures specifically tailored to ERP data characteristics. Our comparison in Table 4 of token-embedding methods for the EEG Transformer may serve as a starting point for designing an ERP-specific Transformer method. Moreover, the lack of a clear advantage of existing foundation models over fully supervised training suggests that ERP-centric pre-training may also be necessary. Most current EEG foundation models are primarily pre-trained on spontaneous EEG corpora, such as TUEG [80], without incorporating ERP-specific preprocessing steps or pre-training strategies. In contrast, ERP data have time-locked trials and stimulus-specific temporal dynamics. These differences suggest that ERP-specific pre-training datasets and preprocessing pipelines are likely essential for improving downstream ERP classification performance.
@@ 7 Conclusion
In this paper, we conduct a comprehensive benchmark study of two ERP tasks: ERP stimulus-type classification and ERP-based brain disease detection, across 12 datasets. We compare 15 methods, including two manual feature-extraction approaches, ten supervised deep learning methods, and three foundation models with pre-trained weights. In addition, to investigate the most suitable token embedding strategy for Transformer-based ERP classification, we conduct a controlled comparison of three commonly used patch embedding methods, aiming to inform the design of ERP-specific Transformer architectures. Based on the results from 36 evaluations, we summarize four key takeaways from this benchmark study. First, deep learning methods consistently outperform manual feature extraction approaches on ERP tasks. Second, existing EEG foundation models do not exhibit clear performance advantages over supervised deep learning models trained from scratch. Third, EEGConformer achieves the most competitive overall performance among the evaluated methods for ERP classification. Fourth, uni-variate patch embedding demonstrates the strongest performance among the examined embedding strategies for ERP-specific Transformers. We hope that this benchmark study serves as a useful reference for method selection in ERP analysis and inspires future research on designing more effective ERP-oriented models for advanced ERP analysis and representation learning.

### v2

@@ 5.2 ERP Stimulus Classification Results
| Method | Stimulus Classification | Disease Detection | Overall Average Rank | ||||||||||
| CESCA- AODD | CESCA- VODD | CESCA- FLANKER | mTBI- ODD | NSERP- MSIT | NSERP- ODD | PD- SIM | PD- ODD | ADHD- WMRI | SCPD | RLPD | AOPD | ||
| EEGConformer | 6.00 | 1.00 | 3.00 | 1.00 | 2.00 | 1.00 | 13.00 | 9.00 | 1.00 | 2.00 | 5.00 | 2.00 | 3.83 |
| ModernTCN | 7.00 | 5.00 | 2.00 | 2.00 | 5.00 | 5.00 | 2.00 | 2.00 | 7.00 | 7.00 | 11.00 | 13.00 | 5.67 |
| TimesNet | 5.00 | 6.00 | 11.00 | 7.00 | 6.00 | 6.00 | 12.00 | 8.00 | 2.00 | 4.00 | 4.00 | 5.00 | 6.33 |
| Medformer | 1.00 | 9.00 | 8.00 | 8.00 | 4.00 | 9.00 | 5.00 | 6.00 | 4.00 | 11.00 | 8.00 | 6.00 | 6.58 |
| CBraMod | 8.00 | 8.00 | 4.00 | 5.00 | 1.00 | 2.00 | 9.00 | 11.00 | 11.00 | 8.00 | 10.00 | 3.00 | 6.67 |
| LaBraM | 2.00 | 11.00 | 9.00 | 10.00 | 9.00 | 8.00 | 10.00 | 5.00 | 5.00 | 3.00 | 1.00 | 8.00 | 6.75 |
| MedGNN | 4.00 | 4.00 | 1.00 | 4.00 | 3.00 | 4.00 | 15.00 | 14.00 | 10.00 | 10.00 | 6.00 | 7.00 | 6.83 |
| PatchTST | 3.00 | 2.00 | 6.00 | 6.00 | 8.00 | 7.00 | 3.00 | 1.00 | 9.00 | 13.00 | 12.00 | 15.00 | 7.08 |
| TCN | 10.00 | 3.00 | 10.00 | 9.00 | 11.00 | 3.00 | 7.00 | 13.00 | 8.00 | 5.00 | 3.00 | 4.00 | 7.17 |
| iTransformer | 9.00 | 7.00 | 7.00 | 3.00 | 7.00 | 11.00 | 1.00 | 3.00 | 6.00 | 9.00 | 14.00 | 14.00 | 7.58 |
| EEGInception | 15.00 | 12.00 | 12.00 | 12.00 | 12.00 | 12.00 | 14.00 | 15.00 | 3.00 | 1.00 | 2.00 | 1.00 | 9.25 |
| ERP Features | 11.00 | 10.00 | 5.00 | 11.00 | 10.00 | 10.00 | 6.00 | 4.00 | 12.00 | 15.00 | 15.00 | 11.00 | 10.00 |
| BIOT | 14.00 | 13.00 | 14.00 | 13.00 | 13.00 | 13.00 | 11.00 | 12.00 | 14.00 | 6.00 | 7.00 | 10.00 | 11.67 |
| EEG Features | 12.00 | 14.00 | 13.00 | 14.00 | 14.00 | 14.00 | 8.00 | 7.00 | 13.00 | 14.00 | 13.00 | 9.00 | 12.08 |
| EEGNet | 13.00 | 15.00 | 15.00 | 15.00 | 15.00 | 15.00 | 4.00 | 10.00 | 15.00 | 12.00 | 9.00 | 12.00 | 12.50 |
@@ 5.3 ERP-Based Disease Detection Results
We then investigate ERP-based disease detection performance, with results summarized in Tables 4. All experiments are conducted under a subject-independent evaluation setup, aiming to identify disease-specific ERP features that generalize across subjects for real-world disease detection applications. Across the six datasets (PD-SIM, PD-ODD, ADHD-WMRI, SCPD, RLPD, and AOPD), the best method achieves F1 scores of 66.79%, 67.37%, 66.43%, 74.33%, 73.87%, and 69.26%, respectively. Overall, no single ERP paradigm consistently demonstrates a clear advantage for disease detection. Notably, the three smaller Parkinson’s disease datasets, SCPD, RLPD, and AOPD, achieve relatively higher performance than the two larger datasets, PD-SIM and PD-ODD. This trend may be attributed to the more balanced data distributions in the smaller datasets, which contain exactly equal numbers of Parkinson’s disease and healthy control subjects. For method comparison, top-tier performance is primarily achieved by deep learning models trained from scratch and by models fine-tuned from foundation model checkpoints. Two methods based on manually extracted features never achieve top-tier results. Besides, unlike ERP stimulus-type classification in the previous subsection, ERP-specific features do not consistently outperform generic EEG features in disease detection. In several datasets, ERP-specific features even perform worse, suggesting that the disease-related information in these datasets may be less tightly aligned with canonical ERP components.
@@ 5.4 Overall Performance Analysis
To provide a comprehensive comparison of methods across multiple tasks and evaluation metrics, we further analyze their overall performance across 12 datasets. For each dataset, we first compute a dataset-level score for method by averaging its Accuracy, F1 score, and AUROC:
where denotes the dataset-level score of method on dataset . Based on these scores, all methods are ranked within each dataset. Let denote the rank of method on dataset . The overall average rank of method across all datasets is then computed as
where is the number of datasets and is the number of compared methods. Figure 4 presents the overall average rank of each method across 12 datasets. Table 5 provides the detailed ranks of each method on each dataset as complementary information. Each value in the table (except for the last column) corresponds to , while the last column reports the overall average rank .
@@
We observe that supervised deep learning models trained from scratch consistently outperform other methods in overall performance. Among all evaluated approaches, EEGConformer achieves the lowest average rank of 3.83, indicating the most stable and competitive performance across diverse tasks and datasets. Following EEGConformer, ModernTCN, TimesNet, Medformer, and CBraMod constitute the top-tier group. Notably, only CBraMod falls within the foundation model category, whereas the remaining methods are trained from scratch. This suggests that existing EEG foundation models do not demonstrate clear performance advantages over supervised deep learning models trained from scratch. In addition, manual feature-based methods consistently rank in the lowest performance tier, while most deep learning approaches easily outperform them. Between the two handcrafted feature methods, ERP-specific features perform relatively better than generic EEG features. These results highlight the limitations of handcrafted statistical and frequency-domain features, which struggle to maintain robust performance across diverse ERP paradigms and disease conditions, particularly under subject-independent evaluation. This suggests that, beyond interpretability or neuroscience-oriented analysis, manual feature extraction offers limited advantages in real-world applications where end-to-end performance is the primary objective.
| Method | Mean F1 | 95% CI | -value |
| LaBraM | 0.0142 | [-0.0005, 0.0290] | 0.0585 |
| ModernTCN | 0.0194 | [0.0052, 0.0335] | 0.0167 |
| iTransformer | 0.0239 | [0.0086, 0.0391] | 0.0083 |
| PatchTST | 0.0252 | [0.0101, 0.0404] | 0.0059 |
| TimesNet | 0.0171 | [0.0071, 0.0271] | 0.0058 |
| Medformer | 0.0194 | [0.0083, 0.0304] | 0.0051 |
| CBraMod | 0.0227 | [0.0111, 0.0342] | 0.0016 |
| TCN | 0.0208 | [0.0113, 0.0302] | |
| MedGNN | 0.0318 | [0.0192, 0.0444] | |
| EEGInception | 0.0507 | [0.0334, 0.0680] | |
| ERP Features | 0.0572 | [0.0404, 0.0740] | |
| EEGNet | 0.1344 | [0.1018, 0.1669] | |
| BIOT | 0.0921 | [0.0697, 0.1144] | |
| EEG Features | 0.1212 | [0.0931, 0.1492] |
| Multi-Variate (25) | Multi-Variate (50) | Multi-Variate (100) | Uni-Variate (25) | Uni-Variate (50) | Uni-Variate (100) | Whole-Variate (—) | |
| CESCA-AODD | 53.260.70 | 53.410.59 | 53.150.69 | 53.910.97 | 53.550.72 | 54.020.73 | 53.350.81 |
| CESCA-VODD | 67.001.62 | 66.101.98 | 67.271.33 | 67.471.83 | 68.271.68 | 67.711.77 | 65.661.33 |
| CESCA-FLANKER | 63.210.84 | 63.760.99 | 63.380.83 | 63.730.87 | 63.871.13 | 63.810.82 | 63.351.39 |
| mTBI-ODD | 63.411.62 | 63.611.61 | 63.181.79 | 63.881.61 | 64.151.91 | 64.511.83 | 63.981.67 |
| NSERP-MSIT | 37.342.21 | 36.872.00 | 36.622.42 | 37.542.31 | 36.652.51 | 36.872.85 | 35.692.37 |
| NSERP-ODD | 65.362.32 | 64.881.77 | 63.622.21 | 65.211.72 | 63.962.55 | 63.592.61 | 62.392.69 |
| PD-SIM | 58.165.05 | 63.462.14 | 64.331.72 | 64.414.28 | 66.592.27 | 68.411.73 | 67.142.55 |
| PD-ODD | 62.453.19 | 63.032.16 | 64.172.08 | 68.030.97 | 68.851.30 | 68.500.44 | 66.131.58 |
| ADHD-WMRI | 63.775.06 | 62.283.82 | 61.302.88 | 61.033.90 | 60.173.94 | 60.533.15 | 60.843.37 |
| SCPD | 67.846.87 | 65.906.48 | 62.187.83 | 65.965.40 | 68.055.66 | 67.915.04 | 65.445.03 |
| RLPD | 64.474.74 | 61.195.68 | 62.514.36 | 62.565.18 | 61.475.46 | 60.734.92 | 59.485.00 |
| AOPD | 65.296.78 | 63.497.32 | 59.964.39 | 58.704.80 | 59.497.46 | 59.436.76 | 59.196.45 |
5.5 Statistical Significance Analysis
We first performed a repeated-measures ANOVA across all compared methods using F1 scores. Since our benchmark adopts a Monte Carlo subject-independent evaluation setup, different seeds correspond to different subject partitions for the train/validation/test split. Therefore, each dataset-seed pair is treated as a repeated evaluation block, with method as the within-block factor. When the ANOVA indicated significant differences, we conducted post-hoc pairwise comparisons between the best-ranked method, EEGConformer, and all other methods using paired -tests with Holm correction.
The repeated-measures ANOVA revealed significant performance differences among the compared methods across the 12 datasets (). Therefore, the null hypothesis that all methods achieve equivalent performance is rejected, and post-hoc paired -tests are conducted to assess pairwise statistical significance. The results are presented in Table 6, which reports the mean F1 difference (F1 = EEGConformer method), the corresponding 95% confidence interval, and the Holm-corrected -value for each pairwise comparison. EEGConformer significantly outperforms most compared methods, while LaBraM shows only a borderline difference, as its confidence interval slightly overlaps zero. Overall, these results provide strong statistical evidence supporting the superiority of EEGConformer for ERP classification, consistent with the average ranking results in Table 5.
5.6 Embedding Method Comparison for Transformer
@@
We compare three commonly used token embedding methods in existing EEG Transformer models to examine their suitability for ERP data classification. Patch lengths of 25, 50, and 100 are used for multi-variate and uni-variate embeddings to eliminate performance fluctuation caused by patch length. All other components, including the self-attention modules, feedforward networks, and layer normalization layers, are kept identical across different embedding methods.
@@
All results are reported in terms of F1 score in Table 7. The uni-variate embedding method achieves the best performance on 8 out of 12 datasets, while the multi-variate embedding method achieves the best results on the remaining 4 datasets. In contrast, the whole-variate embedding method does not achieve top performance on any dataset. These results demonstrate the advantage of uni-variate token embedding for ERP classification and suggest promising start directions for designing ERP-specific Transformer backbones. Moreover, we observe that finer self-attention with shorter patch lengths yields advantages only in the multi-variate token embedding setting, where all 4 best-performing cases occur with a patch length of 25. In contrast, uni-variate token embedding does not consistently achieve the best performance at a specific patch length, and the optimal patch length varies across datasets. This observation suggests that appropriate patch-length selection remains important when using uni-variate token embeddings.
@@ 6 Discussion
Overall, performance on both stimulus-type classification and brain disease detection remains relatively limited for real-world deployment. This limitation may stem from a combination of dataset characteristics and methodological choices. Although the number of trials in existing ERP datasets is often sufficient for traditional manual feature-based analysis, it may still be inadequate for training data-hungry deep learning models. In addition, dataset quality factors, such as noise contamination and class imbalance, can further constrain achievable performance, as discussed in section 5.3 by comparing Parkinson’s disease detection performance across different datasets.
From a method perspective, EEGConformer, a model proposed several years ago, achieves the best average ranking in overall performance, outperforming more recent approaches, including foundation models. This is notable given that foundation models typically contain more parameters and require substantially longer training time, even when fine-tuned from pre-trained checkpoints. This observation suggests that recent advances in general EEG representation learning do not directly translate to the ERP sub-domain, highlighting the need for backbone architectures specifically tailored to ERP data characteristics. The comparison of embedding methods (Table 7) for the Transformer provides a useful starting point for designing ERP-specific Transformer architectures. The results suggest that uni-variate embedding may be a better choice for token embedding, although the optimal patch length remains challenging to determine. Future work on exploring multi-scale learning strategies may be a solution, similar to those used in Medformer [73].
Moreover, the lack of a clear advantage of existing foundation models over fully supervised training suggests that ERP-centric pre-training may also be necessary. Most current EEG foundation models are primarily pre-trained on spontaneous EEG corpora. For example, BIOT was pre-trained on the resting-state EEG dataset PREST [77]; LaBraM was pre-trained on 16 datasets totaling 2,500 hours, among which only 4 datasets (81.16 hours) contain ERP recordings; and CBraMod was trained on TUEG [79], which consists entirely of spontaneous EEG data. Furthermore, none of these models incorporates ERP-specific preprocessing procedures or pre-training strategies. However, ERP data are organized as time-locked trials with stimulus-specific temporal dynamics, which differ fundamentally from spontaneous EEG. These domain gaps suggest that ERP-oriented datasets, preprocessing pipelines, and pre-training methods are likely crucial for improving downstream ERP classification performance.
@@ 7 Conclusion
In this paper, we conduct a comprehensive benchmark study of two ERP tasks: ERP stimulus-type classification and ERP-based brain disease detection, across 12 datasets. We compare 15 methods, including two manual feature-extraction approaches, ten supervised deep learning methods, and three EEG foundation models with pre-trained weights. In addition, to investigate the most suitable token embedding strategy for Transformer-based ERP classification, we conduct a controlled comparison of three commonly used patch embedding methods, aiming to inform the design of ERP-specific Transformer architectures. Based on the results from 36 evaluations, we summarize four key takeaways from this benchmark study. First, modern deep learning methods consistently outperform manual feature extraction approaches on ERP tasks. Second, existing EEG foundation models do not exhibit clear performance advantages over supervised deep learning models trained from scratch. Third, EEGConformer achieves the most competitive overall performance among the evaluated methods for ERP classification. Fourth, uni-variate patch embedding demonstrates the strongest performance among the examined embedding strategies for ERP-specific Transformers. We hope that this benchmark study serves as a useful reference for method selection in ERP analysis and inspires future research on designing more effective ERP-oriented models for advanced ERP analysis and representation learning.